When John F. Kennedy was Sworn in as President, he was 1 year older than Teddy Roosevelt was when Roosevelt took the office. If the sum of their when each became President was 85, how wold was Kennedy when he was sworn in?
Answer by checkley75(3666)

X+(X+1)=85
2X+1=85
2X=85-1
2X=84
X=84/2
X=42 ANSWER FOR THE AGE OF KENNEDY.
ROOSEVELT WAS 42+1=43
PROOF
42+43=85
85=85